The Richibouctou Inn is trying to determine its break-even point. The inn has 70 rooms available that are rented at $50 a night. Operating costs are as follows:

Salaries $10,680 per month Maintenance $400 per month
Utilities $1,950 per month Housekeeping service $4 per room
Depreciation $1,010 per month Other costs $26 per room

Determine the inn’s break-even point in the number of rented rooms per month and dollars.

Break-even point rooms
Break-even point $

If the inn plans on renting 30 rooms per day (assuming a 30-day month), what is the monthly margin of safety in dollars and the margin of safety ratio?

Monthly margin of safety $
Margin of safety ratio %
Answer #1

a)

Fixed Cost = 10,680 + 1,950 + 1,010 + 400 

                  = 14,040

Break-Even Point = FC / CMpr

                             = 14,040 / [50 - (26 + 4)]

Room = 702

Break-Even Point = 702 * 50

                             = $35,100

b)

Actual Sales = day * room * cost per night

                     = 30 * 30 * 50

                     = 45,000

             BEP = 35,100

             MoS = 9,900

           MoSR = 22% (9,900 / 45,000)

answered by: JAS138
Answer #2

Fixed Cost = 10680+1950+1010+400 = 14040

Contribution margin per room = 50-30 = 20

a) Break even point = 14040/20 = 702 rooms

Break even point = 702*50 = $35100

b) Margin of safety = (30*30)-702 = 198*50 = $9900

Margin of safety (%) = 198/900 = 22%